# Implications for the non-Gaussianity of primordial gravitational waves from pulsar timing arrays

### v1

@@ Abstract
Introduction. Primordial gravitational waves (PGWs), as one of the key predictions of inflationary cosmology, offer an exceptional opportunity to probe the early Universe Starobinsky (1979); Rubakov et al. (1982). The standard inflationary paradigm predicts that PGWs follow a nearly Gaussian distribution due to the linearity of quantum field perturbations. However, certain extensions to the inflationary framework introduce non-Gaussianities in PGWs, manifesting as deviations from Gaussian statistics in correlation functions. Investigating non-Gaussian features provides an opportunity to explore inflationary dynamics beyond the simplest models, offering deeper insights into the fundamental physics underlying the early Universe.
@@
Non-Gaussianities in PGWs arise in various inflationary scenarios, such as extra fields Shiraishi et al. (2011a); Shiraishi (2012); Agrawal et al. (2018a); Dimastrogiovanni et al. (2019); Goon et al. (2019); Agrawal et al. (2018b); Gorji et al. (2020); Iacconi et al. (2020), non-attractor phase for tensor fluctuations Ozsoy et al. (2019), massive gravity theory Domènech et al. (2017); Fujita et al. (2020), generalized G-inflation Gao et al. (2011, 2013); Choudhury and Pal (2015); Huang et al. (2013), non-standard vacuum states Akama et al. (2020); Naskar and Pal (2022); Kanno and Sasaki (2022); Gong et al. (2023), higher-derivative Weyl terms De Luca et al. (2019); Maldacena and Pimentel (2011); Shiraishi et al. (2011b), Chern-Simons interactions Lue et al. (1999); Alexander and Martin (2005); Soda et al. (2011); Shiraishi et al. (2011b); Bartolo and Orlando (2017); Bartolo et al. (2019a); Mylova (2019); Bartolo et al. (2021), and more generally, effective field theory Naskar and Pal (2018, 2020); Bordin and Cabass (2020); Cabass et al. (2022). Constraints on tensor non-Gaussianity can be derived from measurements of the three-point correlation function of the cosmic microwave background (CMB) temperature anisotropies and polarization. Up to now, the theoretical investigation of different non-Gaussianity shapes and their detectability has been conducted. Moreover, some of these shapes have already been evaluated using CMB data Shiraishi and Sekiguchi (2014); Shiraishi et al. (2015, 2018); Akrami et al. (2020); Shiraishi (2019); Philcox and Shiraishi (2024). However, large-scale non-Gaussianity constraints alone are insufficient to distinguish between different inflationary models. With the upcoming launch of space-based GW interferometers, such as LISA Amaro-Seoane et al. (2017), Taiji Ruan et al. (2020), and TianQin Luo et al. (2016), it is crucial to probe small-scale tensor non-Gaussianity to complete the puzzle of full-scale tensor non-Gaussianity.
@@
Usually, tensor non-Gaussianity primarily impacts the PGW power spectrum by introducing nontrivial spectral tilts, which can enhance or suppress the GW across different frequency ranges. Additionally, non-Gaussian corrections may shift the peak frequency of the PGW, pushing it to either higher or lower frequencies. This sensitivity to tensor non-Gaussianity offers a pathway to distinguish between various inflationary models and to probe the fundamental dynamics of inflation. Recently, multiple pulsar timing array (PTA) collaborations—such as the Chinese PTA (CPTA) Xu et al. (2023), the European PTA (EPTA) in collaboration with the Indian PTA (InPTA) Antoniadis et al. (2023a, b), the Parkes PTA (PPTA) Zic et al. (2023); Reardon et al. (2023), the North American Nanohertz Observatory for GWs (NANOGrav) Agazie et al. (2023a, b), and the MeerKAT PTA (MPTA) Miles et al. (2024a, b) have independently unveiled robust evidence suggesting the presence of a stochastic gravitational-wave background (SGWB). If the observed signal can be attributed to PGWs, its power spectrum may exhibit observable effects induced by non-Gaussianity. PTA data analysis can therefore constrain non-Gaussianity parameters associated with PGWs.
@@
In this letter, we consider the impact of tensor non-Gaussianity on the power spectrum of PGWs. As the first step, we only focus on the local-type non-Gaussianities, parameterized by a tensor non-Gaussianity parameter $\mathcal{F}_{\mathrm{NL}}$. Using the latest PTA datasets, including the NANOGrav 15-year data release, PPTA DR3, and EPTA DR2, we perform a joint constraints on $\mathcal{F}_{\mathrm{NL}}$. We find $-7.97\lesssim\mathcal{F}_{\mathrm{NL}}\lesssim 7.97$ when primordial tensor perturbation follows a log-normal form. This result establishes a robust framework for testing inflationary models and predictions of the statistical properties of PGWs.
@@
Recent studies suggest that second-order scalar-induced GWs could also account for the PTA signals Inomata et al. (2024); Chen et al. (2020); Liu et al. (2024a); Franciolini et al. (2023); Hosseini Mansoori et al. (2023); Wang et al. (2024); Jin et al. (2023); Liu et al. (2023); Zhu et al. (2024); Yi et al. (2024a); Harigaya et al. (2023); Balaji et al. (2023); Yi et al. (2024b); You et al. (2023); Liu et al. (2024b); Choudhury et al. (2024a, 2023); Domènech et al. (2024); Chen and Liu (2024); Choudhury et al. (2024b); Choudhury and Sami (2025); Chen et al. (2024); Han et al. (2024); Choudhury et al. (2024c). In this context, GWs induced by non-Gaussian curvature perturbations have attracted considerable attention Cai et al. (2019a); Unal (2019); Yuan and Huang (2021); Adshead et al. (2021); Yuan et al. (2023); Zeng et al. (2024), and one can use PTA data to constrain the non-Gaussianity of curvature perturbations on small scales Liu et al. (2024a); Franciolini et al. (2023); Wang et al. (2024); Chang et al. (2024). We emphasize that although scalar-induced GWs also exhibit non-Gaussianity, this non-Gaussianity arises from the curvature perturbations themselves, rather than from the tensor non-Gaussianity inherent to the GWs. Our work serves as an important supplement to the study of scalar non-Gaussianity and will provide valuable insights into exploring the physics of inflation.
@@
Primordial gravitational waves. To investigate the impact of tensor non-Gaussianity on the power spectrum of PGWs, we consider parameterizing the non-Gaussianity of PGWs. A straightforward approach is to mimic the local expansion used in characterizing primordial scalar perturbations. Specially, we assume that the deviation from Gaussianity is small, leading to tensor perturbations in real space expressed as the sum of a Gaussian component and its quadratic term Komatsu and Spergel (2001); Babich et al. (2004); Bartolo et al. (2018)
@@
Combining Eqs. (Implications for the non-Gaussianity of primordial gravitational waves from pulsar timing arrays), (4) and (5), we obtain
@@
where $A$ denotes the amplitude of the power spectrum, $k_{*}$ is the pivot scale corresponding to the peak location, and $\Delta$ characterizes the width of the spectrum.
Methodology and result. To investigate non-Gaussianity of PGWs, we perform a joint analysis of data from three major PTA collaborations: the NANOGrav 15-year dataset Agazie et al. (2023b), PPTA DR3 Zic et al. (2023), and EPTA DR2 Antoniadis et al. (2023a). These long-term observational campaigns represent the culmination of over a decade of precise pulsar timing measurements, with each collaboration independently reporting evidence for a SGWB showing the characteristic Hellings-Downs Hellings and Downs (1983) spatial correlations. The datasets encompass observations from complementary pulsar observations: NANOGrav monitors 68 pulsars over 16.03 years Agazie et al. (2023b), PPTA tracks 32 pulsars for up to 18 years Zic et al. (2023), and EPTA observes 25 pulsars across 24.7 years Antoniadis et al. (2023a). The detection of consistent signals Agazie et al. (2024) across these independent arrays strengthens the case for a GW origin, motivating our joint analysis approach to maximize parameter estimation precision.
@@
Our analysis compares two competing models for the PGW background: a standard Gaussian model ($\mathcal{M}_{\mathrm{G}}$) and a non-Gaussian model ($\mathcal{M}_{\mathrm{NG}}$) incorporating primordial tensor non-Gaussianity. The resulting parameter constraints, displayed in Fig. 2, reveal a significant degeneracy between the amplitude parameter $A$ and the non-Gaussianity parameter $\mathcal{F}_{\mathrm{NL}}$ in the $\mathcal{M}_{\mathrm{NG}}$ analysis, arising from their similar effects on the overall signal strength near the spectrum’s peak. Despite this degeneracy, our joint analysis yields constraints, with the amplitude bounded to $A=0.39^{+0.43}_{-0.27}$ (90% C.L.) and non-Gaussianity parameter constrained to $|\mathcal{F}_{\mathrm{NL}}|\lesssim 7.97$ (90% C.L.). Due to the symmetry in how non-Gaussian corrections enter our model, we find equivalent constraints for positive and negative values, leading to the bilateral bound $-7.97\lesssim\mathcal{F}_{\mathrm{NL}}\lesssim 7.97$, with complete parameter estimates and their uncertainties presented in Table 1.
Conclusion and discussion. Inflation is a cornerstone of modern cosmology, providing insights into the dynamics of the early Universe and aligning well with observational data. However, key questions about the inflationary mechanism remain unresolved, such as the energy scale of inflation and its particle content. The most promising observables for addressing these questions are the power spectrum and bispectrum of primordial perturbations, both in the scalar and tensor sectors. Current data impose stringent constraints on the inflationary scalar sector at CMB scales. In contrast, the tensor sector may provide vital insights into the mass and spin of extra fields and is directly connected to modified gravity. Thus, exploring tensor non-Gaussianity complements the scalar one, helping to test various inflationary models.
In this letter, assuming the signal detected by the PTA collaborations originates from PGWs, we jointly utilize the NANOGrav 15-yr data set, PPTA DR3, and EPTA DR2 to constrain the PGWs with non-Gaussianity. For the first time, we place a limit on the non-Gaussianity parameter $-7.97\lesssim\mathcal{F}_{\mathrm{NL}}\lesssim 7.97$ for a log-normal power spectrum of PGWs. Moreover, as illustrated in Figure 1, the energy density spectrum of PGWs can typically be extended to the frequency range detectable by space-based GW observatories. Consequently, the combined multi-band observations from PTAs and the upcoming space-borne GW detectors, including LISA Amaro-Seoane et al. (2017), Taiji Ruan et al. (2020), and TianQin Luo et al. (2016), will offer a complementary approach to studying tensor non-Gaussianity.

### v2

@@ Abstract
Primordial gravitational waves (PGWs), as one of the key predictions of inflationary cosmology, offer an exceptional opportunity to probe the early Universe Starobinsky (1979); Rubakov et al. (1982). The standard inflationary paradigm predicts that PGWs follow a nearly Gaussian distribution due to the linearity of quantum field perturbations. However, certain extensions to the inflationary framework introduce non-Gaussianities in PGWs, manifesting as deviations from Gaussian statistics in correlation functions. Investigating non-Gaussian features provides an opportunity to explore inflationary dynamics beyond the simplest models, offering deeper insights into the fundamental physics underlying the early Universe.
@@
Non-Gaussianities in PGWs arise in various inflationary scenarios, such as extra fields Shiraishi et al. (2011a); Shiraishi (2012); Agrawal et al. (2018a); Dimastrogiovanni et al. (2019); Goon et al. (2019); Agrawal et al. (2018b); Gorji et al. (2020); Iacconi et al. (2020), non-attractor phase for tensor fluctuations Ozsoy et al. (2019); Mylova et al. (2018), massive gravity theory Domènech et al. (2017); Fujita et al. (2020), generalized G-inflation Gao et al. (2011, 2013); Choudhury and Pal (2015); Huang et al. (2013), non-standard vacuum states Akama et al. (2020); Naskar and Pal (2022); Kanno and Sasaki (2022); Gong et al. (2023), higher-derivative Weyl terms De Luca et al. (2019); Maldacena and Pimentel (2011); Shiraishi et al. (2011b), Chern-Simons interactions Lue et al. (1999); Alexander and Martin (2005); Soda et al. (2011); Shiraishi et al. (2011b); Bartolo and Orlando (2017); Bartolo et al. (2019a); Mylova (2019); Bartolo et al. (2021), and more generally, effective field theory Naskar and Pal (2018, 2020); Bordin and Cabass (2020); Cabass et al. (2022). Constraints on tensor non-Gaussianity can be derived from measurements of the three-point correlation function of the cosmic microwave background (CMB) temperature anisotropies and polarization. Up to now, the theoretical investigation of different non-Gaussianity shapes and their detectability has been conducted. Moreover, some of these shapes have already been evaluated using CMB data Shiraishi and Sekiguchi (2014); Shiraishi et al. (2015, 2018); Akrami et al. (2020); Shiraishi (2019); Philcox and Shiraishi (2024). However, large-scale non-Gaussianity constraints alone are insufficient to distinguish between different inflationary models. With the upcoming launch of space-based GW interferometers, such as LISA Amaro-Seoane et al. (2017), Taiji Ruan et al. (2020), and TianQin Luo et al. (2016), it is crucial to probe small-scale tensor non-Gaussianity to complete the puzzle of full-scale tensor non-Gaussianity.
@@
Usually, tensor non-Gaussianity primarily impacts the PGW power spectrum by introducing nontrivial spectral tilts, which can enhance or suppress the GW across different frequency ranges. Additionally, non-Gaussian corrections may shift the peak frequency of the PGW, pushing it to either higher or lower frequencies. This sensitivity to tensor non-Gaussianity offers a pathway to distinguish between various inflationary models and to probe the fundamental dynamics of inflation. Recently, multiple pulsar timing array (PTA) collaborations—such as the Chinese PTA (CPTA) Xu et al. (2023), the European PTA (EPTA) in collaboration with the Indian PTA (InPTA) Antoniadis et al. (2023a, b), the Parkes PTA (PPTA) Zic et al. (2023); Reardon et al. (2023), the North American Nanohertz Observatory for GWs (NANOGrav) Agazie et al. (2023a, b), and the MeerKAT PTA (MPTA) Miles et al. (2024a, b) have independently unveiled robust evidence suggesting the presence of a stochastic gravitational-wave background (SGWB). If the observed signal can be attributed to PGWs Ye et al. (2024); Jiang et al. (2024); Cai et al. (2021); Fu et al. (2024); Vagnozzi (2023); Choudhury (2024); Ben-Dayan et al. (2023); Oikonomou (2023a, b), its power spectrum may exhibit observable effects induced by non-Gaussianity. PTA data analysis can therefore constrain non-Gaussianity parameters associated with PGWs.
@@
In this paper, we consider the impact of tensor non-Gaussianity on the power spectrum of PGWs. As the first step, we only focus on the local-type non-Gaussianities, parameterized by a tensor non-Gaussianity parameter $\mathcal{F}_{\mathrm{NL}}$. Using the latest PTA datasets, including the NANOGrav 15-year data release, PPTA DR3, and EPTA DR2, we perform a joint constraints on $\mathcal{F}_{\mathrm{NL}}$. We find $-7.97\lesssim\mathcal{F}_{\mathrm{NL}}\lesssim 7.97$ when primordial tensor perturbation follows a log-normal form. This result establishes a robust framework for testing inflationary models and predictions of the statistical properties of PGWs. The choice of a log-normal spectrum is motivated by its ability to model localized features in the PGW power spectrum, which can arise in various inflationary models that induce non-Gaussianity. Several mechanisms can lead to enhancements in the PGW spectrum at specific scales: axion-gauge field inflation Agrawal et al. (2018a, b), higher-derivative Weyl terms De Luca et al. (2019); Maldacena and Pimentel (2011); Shiraishi et al. (2011b), and non-attractor phases of tensor modes Ozsoy et al. (2019); Mylova et al. (2018). In these scenarios, the PGW spectrum often exhibits peaked features, which can be well approximated by a log-normal shape over a relevant frequency range.
@@
Recent studies suggest that second-order scalar-induced GWs could also account for the PTA signals Inomata et al. (2024); Chen et al. (2020); Liu et al. (2024a); Franciolini et al. (2023); Hosseini Mansoori et al. (2023); Wang et al. (2024); Jin et al. (2023); Liu et al. (2023); Zhu et al. (2024); Yi et al. (2024a); Harigaya et al. (2023); Balaji et al. (2023); Yi et al. (2024b); You et al. (2023); Liu et al. (2024b); Choudhury et al. (2024a, 2023); Domènech et al. (2024); Lewicki et al. (2024); Chen and Liu (2024); Choudhury et al. (2024b); Choudhury and Sami (2025); Chen et al. (2024); Han et al. (2024); Choudhury et al. (2024c). In this context, GWs induced by non-Gaussian curvature perturbations have attracted considerable attention Cai et al. (2019a); Unal (2019); Yuan and Huang (2021); Adshead et al. (2021); Yuan et al. (2023); Inui et al. (2024); Zeng et al. (2024), and one can use PTA data to constrain the non-Gaussianity of curvature perturbations on small scales Liu et al. (2024a); Franciolini et al. (2023); Wang et al. (2024); Chang et al. (2024). We emphasize that although scalar-induced GWs also exhibit non-Gaussianity, this non-Gaussianity arises from the curvature perturbations themselves, rather than from the tensor non-Gaussianity inherent to the GWs. Our work serves as an important supplement to the study of scalar non-Gaussianity and will provide valuable insights into exploring the physics of inflation.
II Primordial gravitational waves.
@@
To investigate the impact of tensor non-Gaussianity on the power spectrum of PGWs, we consider parameterizing the non-Gaussianity of PGWs. A straightforward approach is to mimic the local expansion used in characterizing primordial scalar perturbations. Specially, we assume that the deviation from Gaussianity is small, leading to tensor perturbations in real space expressed as the sum of a Gaussian component and its quadratic term Komatsu and Spergel (2001); Babich et al. (2004); Bartolo et al. (2018)
@@
Combining Eqs. (II), (4) and (5), we obtain
@@
where $A$ denotes the amplitude of the power spectrum, $k_{*}$ is the pivot scale corresponding to the peak location, and $\Delta$ characterizes the width of the spectrum. The form of the power spectrum is crucial for understanding the observational consequences of tensor non-Gaussianity. Instead of assuming a simple power-law spectrum, we adopt a log-normal parametrization, which is a useful phenomenological approach to describe localized features in the PGW spectrum. This choice is motivated by several inflationary models that predict localized tensor power enhancements. For example, axion-gauge field interactions Agrawal et al. (2018a, b) can lead to sharp peaks in the PGW spectrum. Similarly, non-attractor phases of inflation Ozsoy et al. (2019); Mylova et al. (2018) can generate scale-dependent enhancements in tensor modes. The log-normal function provides a smooth, well-behaved approximation to such features, making it an ideal choice for our analysis.
III Methodology and result.
To investigate non-Gaussianity of PGWs, we perform a joint analysis of data from three major PTA collaborations: the NANOGrav 15-year dataset Agazie et al. (2023b), PPTA DR3 Zic et al. (2023), and EPTA DR2 Antoniadis et al. (2023a). These long-term observational campaigns represent the culmination of over a decade of precise pulsar timing measurements, with each collaboration independently reporting evidence for a SGWB showing the characteristic Hellings-Downs Hellings and Downs (1983) spatial correlations. The datasets encompass observations from complementary pulsar observations: NANOGrav monitors 68 pulsars over 16.03 years Agazie et al. (2023b), PPTA tracks 32 pulsars for up to 18 years Zic et al. (2023), and EPTA observes 25 pulsars across 24.7 years Antoniadis et al. (2023a). The detection of consistent signals Agazie et al. (2024) across these independent arrays strengthens the case for a GW origin, motivating our joint analysis approach to maximize parameter estimation precision.
@@
Our analysis compares two competing models for the PGW background: a standard Gaussian model ($\mathcal{M}_{\mathrm{G}}$) and a non-Gaussian model ($\mathcal{M}_{\mathrm{NG}}$) incorporating primordial tensor non-Gaussianity. In both cases, we assume a log-normal power spectrum, which serves as a phenomenological template for localized tensor power enhancements. This choice is motivated by theoretical models of inflation that predict peaked features in the PGW spectrum, as discussed in Section II. While alternative parametrizations exist, the log-normal form provides a mathematically convenient and physically meaningful approximation for such enhancements. The resulting parameter constraints, displayed in Fig. 2, reveal a significant degeneracy between the amplitude parameter $A$ and the non-Gaussianity parameter $\mathcal{F}_{\mathrm{NL}}$ in the $\mathcal{M}_{\mathrm{NG}}$ analysis, arising from their similar effects on the overall signal strength near the spectrum’s peak. Despite this degeneracy, our joint analysis yields constraints, with the amplitude bounded to $A=0.39^{+0.43}_{-0.27}$ (90% C.L.) and non-Gaussianity parameter constrained to $|\mathcal{F}_{\mathrm{NL}}|\lesssim 7.97$ (90% C.L.). Due to the symmetry in how non-Gaussian corrections enter our model, we find equivalent constraints for positive and negative values, leading to the bilateral bound $-7.97\lesssim\mathcal{F}_{\mathrm{NL}}\lesssim 7.97$, with complete parameter estimates and their uncertainties presented in Table 1.
IV Conclusion and discussion.
Inflation is a cornerstone of modern cosmology, providing insights into the dynamics of the early Universe and aligning well with observational data. However, key questions about the inflationary mechanism remain unresolved, such as the energy scale of inflation and its particle content. The most promising observables for addressing these questions are the power spectrum and bispectrum of primordial perturbations, both in the scalar and tensor sectors. Current data impose stringent constraints on the inflationary scalar sector at CMB scales. In contrast, the tensor sector may provide vital insights into the mass and spin of extra fields and is directly connected to modified gravity. Thus, exploring tensor non-Gaussianity complements the scalar one, helping to test various inflationary models.
In this paper, assuming the signal detected by the PTA collaborations originates from PGWs, we jointly utilize the NANOGrav 15-yr data set, PPTA DR3, and EPTA DR2 to constrain the PGWs with non-Gaussianity. For the first time, we place a limit on the non-Gaussianity parameter $-7.97\lesssim\mathcal{F}_{\mathrm{NL}}\lesssim 7.97$ for a log-normal power spectrum of PGWs. Moreover, as illustrated in Figure 1, the energy density spectrum of PGWs can typically be extended to the frequency range detectable by space-based GW observatories. Consequently, the combined multi-band observations from PTAs and the upcoming space-borne GW detectors, including LISA Amaro-Seoane et al. (2017), Taiji Ruan et al. (2020), and TianQin Luo et al. (2016), will offer a complementary approach to studying tensor non-Gaussianity.